OpenStudy (anonymous):

Use the product-to-sum formula to write the given product as a sum or difference 12sin(pi/6)sin(pi/6)

OpenStudy (anonymous):

use the identity \[\sin 2 x=2\sin x \cos x\]

OpenStudy (anonymous):

where \[x=\pi/6\]

OpenStudy (anonymous):

2 sin A sin B=cos(A-B)-cos(A+B)